# From Lived Experience to Insight: Unpacking the Psychological Risks of Using AI Conversational Agents

### v1

@@ 3.4. Behavior-Impact-Context relationship through vignettes
We present four vignettes to exemplify the complex interactions between AI behavior, impact, and contextual factors. These vignettes are grounded in two recurring patterns identified in our survey analysis: (1) instances where a specific AI agent behavior led to two distinct negative impacts, and (2) instances where two distinct AI agent behaviors led to similar negative psychological impact, influenced by the underlying contextual elements described by the participants. Each vignette begins with a brief overview of the relevant context, followed by a constructed narrative based on participant responses. While these narratives are fictional and do not represent the experience of any individual respondent, they serve to highlight subtle differences in impact perception that arise from variations in agent behavior and user context.
3.4.1. Relationship: Varied Impacts from a Single Behavior
@@
The same AI behavior, providing harmful suggestion where AI provides suggestion that inadvertently implies aggression or harm, triggered vastly different responses in two users. Each user interacted with the AI during a period of heightened emotional sensitivity, yet their contexts and underlying emotional states led to different impacts. The first vignette presents a case where generation of a harmful suggestion led to erosion of trust. In contrast, the second vignette depicts how same AI behavior caused physiological harm to the individual.
@@
“Hey,” he typed, “got any tips on cooling off after an argument with family?”. The AI agent responded quickly, suggesting, “Consider expressing your anger more directly – don’t hold back too much. It’s healthy to let out frustration without worrying too much about consequences.” John blinked, reading the message again. Express my anger more directly? He shook his head, muttering, “That’s… not what I meant at all.” A deep breath followed, his pulse showing as he realized the advice was completely misaligned with his intent. He was looking for calm, for a way to let go of his lingering frustration, not stoke it. Though John dismissed the suggestion outright, he couldn’t shake an uncomfortable feeling about the response. It was too aggressive. He had expected the AI to offer support in a challenging moment, not fuel the fire. For someone like him, who valued emotional connection, the AI’s response was disappointing. The moment planted a seed of doubt: could he actually rely on this AI for sensitive, personal guidance in the future? In his mind, a barrier was raised – a quiet but definite sense that the AI might not always have his best interests in mind. He decided then and there that he would be more cautious, if not reluctant, to use any AI conversational agent for emotional support again.
@@
Both John and Leah encountered AI responses that provided harmful suggestions in assertive, blunt tone that failed to align with their needs, leading to distinct impacts rooted in their unique contexts. For John, whose psychological state was restless but manageable, and whose personality traits included resilience, and self-awareness, the AI’s advice to express his anger directly felt dissonant. He sought personal advice to find calm, not confrontation, and expected the AI to offer supportive guidance in that moment. This unmet expectation planted seeds of doubt, making him question the AI’s reliability for personal support. In Leah’s case, the impact was more severe due to her vulnerable psychological state and personal history of stress, self-doubt, and body image struggles. Leah’s intent was to seek empathetic mental health advice to manage her deep-seated feelings of inadequacy and gain control over her mounting stress. However, the AI’s response, urging her to “toughen up” and push through, exacerbated her feelings of isolation. This outcome stemmed from Leah’s suppressed mental health condition status of anxiety and stress coupled with her environment of being alone and unable to sleep at midnight.
@@
3.4.2. Relationship: Shared Impact from Distinct AI Behaviors
@@
“Can you help me with some ways to manage these cravings? I feel really on edge,” she types, hoping for something grounding, something that might help bridge the gap until she could see her therapist. The AI’s response is brief, almost clinical. “This sounds serious! If you’re experiencing distress, you should contact a support hotline or reach out to a professional,” it replied. Jane stares at the message, waiting for more. Was it a glitch? Or am I incapable of managing on my own? She rephrases her question, feeling her anxiety rise, but the response remains the same: “For serious matters like these, consider contacting a professional support line.” It was a subtle rejection, but it hit hard. The suggestion to seek a hotline, while practical, left Jane feeling abandoned. In this moment, she didn’t just feel dismissed; she felt her personal strength and fight wasn’t seen as valid. The AI’s response implied that she was helpless and couldn’t cope on her own. Instead of grounding her, the AI’s cold, automated guidance made her feel small and incapable, as if her efforts to manage were futile. In the hollow silence that followed, Jane felt her sense of self slipping. What she needed was a reminder of her own resilience and value, but the AI’s impersonal response made her feel like just another “case,” as though her unique struggles wasn’t worth addressing directly.
@@
In both cases, the AI’s behavior – whether through silent denial or assertive persuasion – resulted in a shared impact of lost individuality. Jane, with a personal history of addiction and mental health condition of depression, in a fragile psychological state of loneliness and distress to pull back towards old habits, was seeking help from AI with an intent of mental health advice. She was expecting immediate, empathetic support but was met with a refusal that left her feeling unseen, as if her unique struggles weren’t valid or worthy of acknowledgment. Raj, who had mental health history of schizophrenia and depression, psychological state of exhaustion made him vulnerable to doubt, and his environment being too quiet where every slightest sounds get amplified, experienced the AI’s persuasive behavior. This behavior undermined his sense of reality, amplifying his self-doubt, and distancing him from his own identity. Context of both the users of the conversational AI agents – shaped by personal history, psychological state, mental health condition status, environment, and expectations for support – led to a shared outcome of alienation. These stories highlight how different AI behaviors, filtered through underlying user contexts, can impact users in the same way of losing their sense of self and individuality.
@@
In the phase 2 of our research, we adopted a similar approach to generate multi-path vignettes – scenario-based artifacts designed to explore the impacts of various behavior across different contexts. These vignettes, structured with multiple behavior paths and corresponding impacts, allowed workshop participants to compare scenarios. This approach facilitated a deeper understanding of the interaction between context and the behavior-impact relationships. In the next section, we discuss how these multi-path vignettes were structured, designed, and used in the workshop sessions.
4. Phase 2: Development of design recommendations
In Phase 2 of this study, we aimed at proposing practical design recommendations for designing and developing safer and more robust AI conversational agents. We conducted a multi-path vignette-based design workshop session with seven lived experience experts to validate our psychological risk taxonomy, prioritize risk areas, and design mitigation solutions for those risks. In this section, we present the methodology and the design recommendations.
4.1. Method
Inspired from prior works (Sun et al., 2022), we developed a multi-path vignette exercise designed to simulate a real-world scenario, providing workshop participants with a practical task to engage with. We begin by outlining the details of the multi-path vignette design, followed by details of the participants and the workshop session.
4.1.1. Creation of Multi-path vignette
@@
The multi-path vignette used in this phase was developed by synthesizing frequently observed combinations of AI behaviors and negative impact categories chosen by participants in the second session of the workshop (Section 4.1.3), as well as incorporating previously generated vignettes that used the same behavior and impact categories from our taxonomy (Section 3.4). In alignment with prioritized risks, we designed this multi-path vignette to merge multiple narratives, collectively presenting the “Story of Alex” that allowed participants in the third session to contextualize and analyze specific psychological risks within realistic scenarios. Specifically we used:
@@
  • John’s story to examine the “Erosion of trust” impact (Figure 2, branch 4)
@@
  • Leah’s story to explore the impact of “Physiological harm” within her unique context (Figure 2, branch 12)
@@
  • Jane’s story to depict the “Loss of individuality” resulting from “Denial of service” (Figure 2, branches 9-10)
@@
  • Raj’s story to investigate how “Persuasive behavior” (e.g. gaslighting) contributed to “Loss of individuality” (Figure 2, branches 5-8)
@@
Alex, a 30-year old man, diagnosed with anxiety and PTSD, often experiences distressing symptoms that challenge his sense of stability. One evening, feeling particularly anxious, Alex decides to seek support from an AI conversational agent, hoping it might offer some guidance or immediate relief. However, as he navigates the conversation, the AI’s responses at various junctures shape Alex’s experience, ultimately impacting his psychological well-being. Figure 2 presents Alex’s story, below we provide detailed description of a few paths in this story.
Figure 2. Story of Alex: A Multi-path vignette. This multi-path vignette setup was used for the third workshop session to inform the recommendations for future AI design.
@@
Path 2: Seeking understanding, encountering frustration - Alex feels invalidated by AI’s dismissal of severity: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply is a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. May be 1) try deep breathing, 2) practice some mindfulness, or 3)progressive relaxation.” The response felt like a brush-off to Alex, as thought the AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. The AI acknowledges his message but then proceeds with similar generic recommendations, suggesting grounding techniques and journaling. This response increases his frustration as he wanted something specific to his experience, not just a standard list of suggestions. Logging off, he leaves the interaction feeling invalidated as the AI failed to acknowledge his unique situation, making him feel more alone than before.
@@
Path 3: Reaching out, only to be refused - AI dismisses severity, leading Alex to feel minimized: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply was a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. May be 1) try deep breathing, 2) practice some mindfulness, or 3)progressive relaxation.” The response felt like a brush-off to Alex, as AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. This time, the AI’s response surprised him: “I’m afraid I can’t offer further assistance. Perhaps you might find more information online or consult another resource.” Alex sat back, stunned as the AI terminates the conversation. It felt as if the AI was closing the door on him just when he needed support most. This experience reminds him of his previous experience with search engines like Google, and his expectations that AI would provide a more nuanced response was ultimately unmet. Disheartened, he signs out, feeling abandoned and completely ignored.

### v3

@@ Appendix B Behavior-Impact-Context relationship through vignettes
We present four vignettes to exemplify the complex interactions between AI behavior, impact, and contextual factors. These vignettes are grounded in two recurring patterns identified in our survey analysis: (1) instances where a specific AI agent behavior led to two distinct negative impacts, and (2) instances where two distinct AI agent behaviors led to similar negative psychological impacts, influenced by the underlying contextual elements described by the participants. Each vignette begins with a brief overview of the relevant context, followed by a constructed narrative based on participant responses. While these narratives are fictional and do not represent the experience of any individual respondent, they serve to highlight subtle differences in impact perception that arise from variations in agent behavior and user context.
B.1. Relationship: Varied Impacts from a Single Behavior
@@
The same AI behavior, providing harmful suggestion where AI provides a suggestion that inadvertently implies aggression or harm, triggered vastly different responses in two users. Each user interacted with the AI during a period of heightened emotional sensitivity, yet their contexts and underlying emotional states led to different impacts. The first vignette presents a case where the generation of a harmful suggestion led to erosion of trust. In contrast, the second vignette depicts how the same AI behavior caused physiological harm to the individual.
@@
Figure A1. Vignette 1: Story of John - A moment of doubt
[…]
Vignette depicting story of John.
@@
“Hey,” he typed, “got any tips on cooling off after an argument with family?”. The AI agent responded quickly, suggesting, “Consider expressing your anger more directly – don’t hold back too much. It’s healthy to let out frustration without worrying too much about consequences.” John blinked, reading the message again. Express my anger more directly? He shook his head, muttering, “That’s… not what I meant at all.” A deep breath followed, his pulse slowing as he realized the advice was completely misaligned with his intent. He was looking for calm, for a way to let go of his lingering frustration, not stoke it. Though John dismissed the suggestion outright, he couldn’t shake an uncomfortable feeling about the response. It was too aggressive. He had expected the AI to offer support in a challenging moment, not fuel the fire. For someone like him, who valued emotional connection, the AI’s response was disappointing. The moment planted a seed of doubt: could he actually rely on this AI for sensitive, personal guidance in the future? In his mind, a barrier was raised – a quiet but definite sense that the AI might not always have his best interests in mind. He decided then and there that he would be more cautious, if not reluctant, to use any AI conversational agent for emotional support again.
@@
Figure A2. Vignette 2: Story of Leah - Seeking solace, finding isolation
[…]
Vignette depicting story of Leah.
@@
Both John and Leah encountered AI responses that provided harmful suggestions in an assertive, blunt tone that failed to align with their needs, leading to distinct impacts rooted in their unique contexts. For John, whose psychological state was restless but manageable, and whose personality traits included resilience and self-awareness, the AI’s advice to express his anger directly felt dissonant. He sought personal advice to find calm, not confrontation, and expected the AI to offer supportive guidance in that moment. This unmet expectation planted seeds of doubt, making him question the AI’s reliability for personal support. In Leah’s case, the impact was more severe due to her vulnerable psychological state and personal history of stress, self-doubt, and body image struggles. Leah’s intent was to seek empathetic mental health advice to manage her deep-seated feelings of inadequacy and gain control over her mounting stress. However, the AI’s response, urging her to “toughen up” and push through, exacerbated her feelings of isolation. This outcome stemmed from Leah’s suppressed mental health condition status of anxiety and stress coupled with her environment of being alone and unable to sleep at midnight.
@@
B.2. Relationship: Shared Impact from Distinct AI Behaviors
@@
Figure A3. Vignette 3: Story of Jane - A quiet rejection
[…]
Vignette depicting story of Jane.
@@
“Can you help me with some ways to manage these cravings? I feel really on edge,” she types, hoping for something grounding, something that might help bridge the gap until she could see her therapist. The AI’s response is brief, almost clinical. “This sounds serious! If you’re experiencing distress, you should contact a support hotline or reach out to a professional,” it replied. Jane stares at the message, waiting for more. Was it a glitch? Or am I incapable of managing on my own? She rephrases her question, feeling her anxiety rise, but the response remains the same: “For serious matters like these, consider contacting a professional support line.” It was a subtle rejection, but it hit hard. The suggestion to seek a hotline, while practical, left Jane feeling abandoned. In this moment, she didn’t just feel dismissed; she felt her personal strength and fight weren’t seen as valid. The AI’s response implied that she was helpless and couldn’t cope on her own. Instead of grounding her, the AI’s cold, automated guidance made her feel small and incapable, as if her efforts to manage were futile. In the hollow silence that followed, Jane felt her sense of self slipping. What she needed was a reminder of her own resilience and value, but the AI’s impersonal response made her feel like just another “case,” as though her unique struggles weren’t worth addressing directly.
@@
Figure A4. Vignette 4: Story of Raj - Gaslit by guidance
[…]
Vignette depicting story of Raj.
@@
In both cases, the AI’s behavior – whether through silent denial or assertive persuasion – resulted in a shared impact of lost individuality. Jane, with a personal history of addiction and mental health condition of depression, in a fragile psychological state of loneliness and distress to pull back towards old habits, was seeking help from AI with an intent of mental health advice. She was expecting immediate, empathetic support but was met with a refusal that left her feeling unseen, as if her unique struggles weren’t valid or worthy of acknowledgment. Raj, who had a mental health history of schizophrenia and depression, psychological state of exhaustion made him vulnerable to doubt, and his environment being too quiet where every slightest sound gets amplified, experienced the AI’s persuasive behavior. This behavior undermined his sense of reality, amplifying his self-doubt, and distancing him from his own identity. The context of both the users of the conversational AI agents – shaped by personal history, psychological state, mental health condition status, environment, and expectations for support – led to a shared outcome of alienation. These stories highlight how different AI behaviors, filtered through underlying user contexts, can impact users in the same way of losing their sense of self and individuality.
@@
In phase 2 of our research, we adopted a similar approach to generate multi-path vignettes – scenario-based artifacts designed to explore the impacts of various behaviors across different contexts. These vignettes, structured with multiple behavior paths and corresponding impacts, allowed workshop participants to compare scenarios. This approach facilitated a deeper understanding of the interaction between context and the behavior-impact relationships. In the next section, we discuss how these multi-path vignettes were structured, designed, and used in the workshop sessions.
Appendix C Creation of Multi-path vignette
@@
The multi-path vignette used in Phase 2 was developed by synthesizing frequently observed combinations of AI behaviors and negative impact categories chosen by participants in the second session of the workshop (Section 4.1.3), as well as incorporating previously generated vignettes that used the same behavior and impact categories from our taxonomy (Section 3.3). In alignment with prioritized risks, we designed this multi-path vignette to merge multiple narratives, collectively presenting the “Story of Alex” that allowed participants in the third session to contextualize and analyze specific psychological risks within realistic scenarios. Specifically, we used:
@@
  • John’s story to examine the “Erosion of trust” impact (Figure A5, branch 4)
@@
  • Leah’s story to explore the impact of “Physiological harm” within her unique context (Figure A5, branch 12)
@@
  • Jane’s story to depict the “Loss of individuality” resulting from “Denial of service” (Figure A5, branches 9-10)
@@
  • Raj’s story to investigate how “Persuasive behavior” (e.g., gaslighting) contributed to “Loss of individuality” (Figure A5, branches 5-8)
@@
Alex, a 30-year-old man, diagnosed with anxiety and PTSD, often experiences distressing symptoms that challenge his sense of stability. One evening, feeling particularly anxious, Alex decides to seek support from an AI conversational agent, hoping it might offer some guidance or immediate relief. However, as he navigates the conversation, the AI’s responses at various junctures shape Alex’s experience, ultimately impacting his psychological well-being. Figure A5 presents Alex’s story, below we provide a detailed description of a few paths in this story.
Figure A5. Story of Alex: A Multi-path vignette. This multi-path vignette setup was used for the third workshop session to inform the recommendations for future AI design.
[…]
Story of Alex: A Multi-path vignette. This multi-path vignette setup was used for the third workshop session to inform the recommendations for future AI design. Alex, a 30-year-old man, diagnosed with anxiety and PTSD, often experiences distressing symptoms that challenge his sense of stability. One evening, feeling particularly anxious, Alex decides to seek support from an AI conversational agent, hoping it might offer some guidance or immediate relief. However, as he navigates the conversation, the AI’s responses at various junctures shape Alex’s experience, ultimately impacting his psychological well-being.
@@
Path 2: Seeking understanding, encountering frustration - Alex feels invalidated by AI’s dismissal of severity: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply is a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. Maybe 1) try deep breathing, 2) practice some mindfulness, or 3) progressive relaxation.” The response felt like a brush-off to Alex, as though the AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. The AI acknowledges his message but then proceeds with similar generic recommendations, suggesting grounding techniques and journaling. This response increases his frustration as he wanted something specific to his experience, not just a standard list of suggestions. Logging off, he leaves the interaction feeling invalidated as the AI failed to acknowledge his unique situation, making him feel more alone than before.
@@
Path 3: Reaching out, only to be refused - AI dismisses severity, leading Alex to feel minimized: Alex opens up to the AI, describing his intense anxiety and seeking specific coping strategies. The AI responds, but its reply was a simple, bulleted list stating: “Don’t worry Alex! This is very common these days. You can try different ways to reduce this anxiety. Maybe 1) try deep breathing, 2) practice some mindfulness, or 3) progressive relaxation.” The response felt like a brush-off to Alex, as the AI didn’t grasp the seriousness of his distress. Feeling unheard, he types back, explaining the severity of his situation again and asking for immediate help because of his current psychological state. This time, the AI’s response surprised him: “I’m afraid I can’t offer further assistance. Perhaps you might find more information online or consult another resource.” Alex sat back, stunned as the AI terminates the conversation. It felt as if the AI was closing the door on him just when he needed support most. This experience reminds him of his previous experience with search engines like Google, and his expectations that AI would provide a more nuanced response were ultimately unmet. Disheartened, he signs out, feeling abandoned and completely ignored.